OpenStudy (anonymous):

-n - 24 = 5 - n SHOW WORK PLEASE

OpenStudy (anonymous):

-1n-24=5-1n +1n +1n 0n-24=5 That does not make any sense because -24 is not equal to 5! Or maybe i did it wrong!